Find three consecutive even integers such that the sum of the first and second equals the sum of the third and -10
Answer by roseo(33)

Let x= the first even integer
x+2 = the second even integer
x+4 = the third even integer (Because we know to get the next even integer we just add 2)
Then the sum of the first two would be x+x+2 and the sum of the third and -10 would be x+4+(-10) or x-6 .
Therefore we can write the equation
x+x+2 = x-6
Combining like terms on the left we have
2x+2 = x-6
Then subtract x from each side
x+2 = -6
Then subtract 2 from each side so we have
x=-8
Then x+2 = -8+2 or -6
and x+4 = -8+4 or -4.
So your three consecutive integers are -8,-6, and-4